OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

use this equation \(\Large y-y_1 = m (x-x_1)\) thats the slope point form

hartnn (hartnn):

here ,x1, y1 = -2,-3 m = slope = -6

OpenStudy (anonymous):

umm im having trouble ding it im really confused

hartnn (hartnn):

plug in the known values! y - (-3) = (-6 ) (x- (2)) can you try to simplify that ?

OpenStudy (anonymous):

is the answer d

OpenStudy (anonymous):

y = -6x + 20

hartnn (hartnn):

nopes, did you try to simplify ?? what is y- (-3) = ... ?

OpenStudy (anonymous):

3

hartnn (hartnn):

you mean y+3 ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

what is x-(-2) = ... ?

OpenStudy (anonymous):

x+2

hartnn (hartnn):

yes y + 3 = -6 (x+2) try to simplify this

OpenStudy (anonymous):

y=-6x-15 ??

hartnn (hartnn):

thats correct! :)

OpenStudy (anonymous):

yayy thank you :))

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

i have another question can you help me?

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (1, -5).

OpenStudy (anonymous):

y = -5 x = 1 y = 1 x = -5

OpenStudy (anonymous):

i think its d

hartnn (hartnn):

horizontal lines have equation of the form y = c where c is the y co-ordinate of any point on the line so, its not D

OpenStudy (anonymous):

so y=-5

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

ok thanks again

hartnn (hartnn):

welcome ^_^